WLCG-OSG-EGEE Ops' Minutes Mon 02 Feb 2009

Summary

Obsoletion of gLite 3.0 tentatively planned for April.
LHCb now require 50GB of shared area capacity on SL5.0 WNs.
Common requirement is for a list of library packages needed for a 32-bit to 64-bit migration.
SARA LFC was down last week, so ATLAS SRMv2 tests at all associated T2s and NIKHEF failed. Request to mark test results somewhat differently.
List of T1 FTM end-points needs consolidating.

Attendance

EGEE

  • Asia Pacific ROC: Jason Shih
  • Central Europe ROC: Malgorzata Krakowian
  • OCC / CERN ROC: John Shade, Antonio Retico, Nick Thackray, Diana Bosio
  • French ROC: Pierre Girard, 0033478930880 (Helene Cordier)
  • German/Swiss ROC: Wen Mei
  • Italian ROC: Absent
  • Northern Europe ROC: Absent
  • Russian ROC: Lev Shamardin
  • South East Europe ROC: Kostas Koumantaros
  • South West Europe ROC: Kai Neuffer
  • UK/Ireland ROC: Jeremy Coles
  • GGUS: Helmut Dres
  • Ticket Police: Maria Dimou
  • Site UNI-KARLSRUHE: Volker Buege
  • Unknown guest: 0081128560187 (Japan)

WLCG

  • WLCG Service Coordination: Harry Renshall

WLCG Tier 1 Sites

  • ASGC: Jason Shih
  • BNL: Absent
  • CERN site: Absent
  • FNAL: Joe Kaiser, Catalin Dumitrescu
  • FZK: Angela Poschlad
  • IN2P3: Pierre Girard
  • INFN: Absent
  • NDGF: Absent
  • PIC: Kai Neuffer
  • RAL: Gareth Smith
  • SARA/NIKHEF: Absent
  • TRIUMF: Absent

LHC Experiments

  • ATLAS: Alessandro di Girolamo
  • LHCb: Roberto Santinelli
  • CMS: absent
  • ALICE: absent

Feedback on Last Week's Minutes

None was given.

EGEE Items

Grid Operator Hand Over on Duty

  Primary Team Secondary Team
From ROC CERN ROC DECH
To ROC Italy ROC France

  • Volker Buege was present to explain why UNI-KARLSRUHE shouldn't be suspended. He explained that their H/W transition had led to considerable problems, but predicted that they’ll be stable again by the end of next week. This satisfied Nick, who agreed that the site shouldn't be suspended.
    Volker also pointed out that CODs shouldn’t open tickets against configuration items (e.g. CE) that are in downtime.
    Does a SE need a working CE for SAM tests? Karlsruhe will bring up BDII and SE services next week, but no CE. Will the SAM tests pass? John said that the BDII and SE tests didn't have a dependency on the CE, but will check what GridView do in terms of calculating availability.
    Update from GridView: if all CEs are scheduled down, overall site will show as scheduled down (even though SE and sBDII will appear as available)
  • Diana: please mask FTS-infosites on fts-t1import.cern.ch which is failing in CIC portal. Helene will follow up with Cyril. Diana will provide GGUS reference [which she did: GGUS:45163, GGUS:44954, GGUS:44635 ].
    Update from Helene: Alarms on this should be closed. A bug has been opened in Savannah and the associated GGUS ticket will be put in the handover ASAP in order to remind next shift not to open tickets on this. Also recorded in https://twiki.cern.ch/twiki/bin/view/EGEE/OperationalUseCasesAndStatus as a reminder for later shifts.
  • As of last Friday, site SDU-LCG2 still hadn’t given any sign of life. After a few more days, they’ll be suspended by Diana.
  • Reminder from Diana for ROCs to register and unregister things in GOCDB as required. Helene reminded people of the CIC portal BDII/GOCDB resource comparator utility, which Diana had never heard of. She’ll send the link to the CIC on duty mailing list, and Kai posted it to the conference whiteboard: https://cic.gridops.org/index.php?section=roc&page=comparator

PPS Reports and Issues

  • Antonio had received news from the gLite release team that the new SCAS (Site Central Authorization Service), complete with documentation, is mature enough for a pilot. LHCb would like a Tier 1 to be involved in the pilot (Lyon & Karlsruhe suggested). A kick-off meeting is planned by Antonio for this week – those interested to contact Antonio, who will add the reference page to the agenda. Pierre: interested, but difficult before end of February. TBD in kickoff.

gLite Release News

  • gLite 3.0 and 3.1 release of VOMS certs for biomed, ATLAS & CMS scheduled for Wednesday. As usual, read the links on the agenda page for details.

EGEE Items From ROC Reports

  • Scratch space discussion (problem also seen at Desy and Bonn). What should sites do? ATLAS need to update their VO cards, but were astonished that 15GB of temporary space was a problem for anyone. However, it would seem that some users do not use the correct temporary location. It was agreed that users causing problems can be killed without mercy (their jobs, that is), but production jobs should be handled by contacting the appropriate VO first.
  • SWE reported some SAM problems over the weekend, and John confirmed that a support call had been opened with the WMS support team.

Obsoletion of gLite 3.0

  • Heads-up from Nick that this will happen at the end of April as long as there are no major objections (to be discussed at TMB). After that, sites will be on their own, with no updates nor support!

WLCG Items

WLCG issues coming from ROC reports

  • SWE would like CMS to update VO requirements on VO-ID card. GGUS ticket has been assigned to VO support CMS.

Upcoming WLCG Service Interventions

  • Consult links on the agenda page.
  • Reminder on agenda concerning scheduled SAM/!GridView DB downtime on Wednesday (2-3 hours)

FTM end points

Current list is on the agenda. Please, information for all T1s needs to be collected and sent to Nick or Maite! Alessandro would like CNAF to update their FTM end-point, because the link provided has been broken for "years".

WLCG Service Coordination

Harry suggested that people review the minutes from the daily coordination meetings, the link to which was kindly supplied by Maria: https://twiki.cern.ch/twiki/bin/view/LCG/WLCGDailyMeetingsWeek090202#Monday

ATLAS Service

  • Alessandro mentioned an ATLAS pre-staging test to take place this week (a few TBytes staged in from tape to disk), and that the order of the clouds used might be different.
  • SARA LFC was down last week, all T2s below couldn’t work, and neither could NIKHEF. ATLAS use catalog of the cloud for SRMv2 tests. If the catalog is down, those tests will fail. Alessandro suggested that this should be fixed by GridView so as not to blame sites for central failures - perhaps by marking result as "test not sent to site"...

ALICE Service

CMS Service

LHCb Service

Roberto summarized the daily meeting, namely that next week, no FEST activity was planned, but that nominal transfers would be carried out.
  • CIC VO card modified: shared area capacity updated to 50GB on SL5.0 WNs.
  • 32-bit to 64 bit migration at PIC: VOs need a list of library packages needed. Who provides it? No one knew for sure, but since this is a general problem, Nick will ask SA3 if they have anything. Kai thanked Roberto for having put useful information on the VO cards - it was a big help to PIC.

OSG Items

Rob was on the line to receive Maria's weekly list of problem tickets.
  • GGUS:44140 user Saul needs pinging, Rob will do this.
  • GGUS:45094 can probably be closed, Maria will check with S. Burke.
  • GGUS:45488 GGUS-OSG interface testing. Will be closed, but need to check why it wasn’t closed automatically. Maria to check with Gunter.

Newly Created Action Items

Assigned to Due date Description State Closed Notify  
NickThackray 2009-02-09 Ask SA3 for a list of library packages needed for 32 to 64-bit migration.

*Update 12th Feb* - There is no list. There is a list of per VO on the VO Cards, we may try and produce a common list. What next?

*Update at the meeting* - VOs will definitely have to maintain a list of the libraries they need, in their VO ID card.
Item closed.

2009-02-27 edit

Assigned to Due date Description State Closed Notify  
AllROCs 2009-02-09 All T1s to check and update list of FTM end-points. To be sent to Nick

12th Feb, TWiki page has been created.

2009-02-12 edit

Review of Open Action Items

Open Action Items

IdSubmitterDescriptionCreationDueAssigned To 

Actions Closed in Last 20 Days

IdSubmitterDescriptionCreationDueAssigned ToClosed 

AOB

None was volunteered.

Next Meeting

The next meeting will be Monday, 09 February 2009 15:00 UTC (16:00 Swiss Mountain Time).

  • Attendees can join from 14:45 UTC (15:45 Swiss local time) onwards.
  • The meeting will start promptly at 15:00 UTC (16:00 Swiss local time).
  • The WLCG section will start at the fixed time of 15:30 UTC (16:30 Swiss local time).
  • To dial in to the conference:
    • Dial +41227676000
    • Enter access code 0148141


These minutes can only be changed by members of:

Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r4 - 2009-02-27 - NickThackray
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    EGEE All webs login

This site is powered by the TWiki collaboration platform Powered by Perl This site is powered by the TWiki collaboration platformCopyright & by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Ask a support question or Send feedback